# Faithfulness-Aware Uncertainty Quantification for Fact-Checking the Output of Retrieval Augmented Generation

### v2

@@ B.3 Automatic Annotation of Claims
As for faithfulness annotation, each claim is assigned one of three categories during the automatic annotation. The categories of faithfulness are “faithful” if the context contains information that directly supports the statement, “unfaithful-contra” if the context contains information that directly contradicts the statement, “unfaithful-neutral” if the context does not contain any information supporting or contradicting the statement. Next, the obtained labels are binarized for experimental evaluation: 1 if labeled faithful and 0 if labeled unfaithful-contra, or unfaithful-neutral. This is done since the unfaithful-contra category is quite small (less than 5% of the data).
Regarding the factuality annotation, the automatic annotation process also assigns one of three categories to each claim. The categories of factuality are “True” if the statement is factually correct according to your knowledge, “False” if the statement is factually incorrect according to your knowledge, “unverifiable” if the factual accuracy of the statement cannot be determined based on available knowledge (should be assessed without referring to the provided context). After that, the annotation for factuality is binarized, namely only verifiable claims are taken (we exclude labels of the unverifiable category), 1 if the claim class is false and 0 if the class is true.
Figure 6 presents the annotation prompt used with GPT-4o-search.
@@ B.4 Manual Enhancement of Automatic Annotation
For manual verification of the automatic annotation, we used reliable sources obtained from Google search results. An initial validation of the automatic labeling was conducted by comparing automatic and manual annotations on randomly selected claims to assess class balance. Specifically, 100 claims were reviewed for Llama 3B and 76 for Falcon 3B. The resulting class distributions are shown in Figure 7(a) and Figure 8(a), respectively. We also present the results of comparison of manual and automatic annotation for faithfulness for both models, see Figure 9(a, b).
Given the difficulty of assessing the false and unverifiable categories in automatic annotations, these cases were prioritized for manual review. Enhanced annotations for the same 100 (Llama 3B) and 76 (Falcon 3B) claims are shown in Figure 7(b) and Figure 8(b). In addition, a full manual re-check of all False and Unverifiable claims was performed, resulting in 359 manually reviewed claims for Llama 3B and 240 for Falcon 3B.
We used six student annotators for this study. Each annotator spent approximately three hours on the annotation task. Instructions were provided orally in the form of informal discussions; no written guidelines were used. All annotators participated voluntarily and were not financially compensated.
To assess annotation consistency, we also conducted an agreement analysis on the 100 Llama 3B claims, each independently reviewed by two annotators (see Table 8 for details). These results suggest that the annotators generally aligned well in both tasks, particularly for factuality, though there is some room for improvement, especially in clarifying ambiguous or borderline cases.
(a) Before manual enhancement of automatic annotation
(b) After manual enhancement of automatic annotation
Figure 7: Balance of classes of factuality annotations for the Llama 3B model. Each matrix is based on 100 randomly selected claims, comparing annotations produced by the model with those from human annotators.
(a) Before manual enhancement of automatic annotation
(b) After manual enhancement of automatic annotation
Figure 8: Balance of classes of factuality annotations for the Falcon 3B model. Each matrix is based on 76 randomly selected claims, comparing annotations produced by the model with those from human annotators.
(a) Llama 3B faithfulness classes
(b) Falcon 3B faithfulness classes
Figure 9: Balance of classes of faithfulness annotations for Llama 3B and Falcon 3B models. The matrices are based on 100 and 76 randomly selected claims, correspondingly, comparing annotations produced by the model with those from human annotators.
@@
Table 8: Inter-annotator agreement for factuality and faithfulness annotations based on 100 claims of Llama 3B. Accuracy measures raw agreement, Cohen’s Kappa adjusts for chance agreement.
Appendix C Additional faithfulness-related results
C.1 Faithfulness Evaluation on Long-Form QA
Table 9 shows the long-form QA dataset results when using faithfulness as the target metric. In this evaluation, all methods remained consistent with those used in the factuality dataset. Among the compared methods, AlignScore achieved the best performance, indicating its effectiveness in approximating the faithfulness within the franq formula.
Method Llama3b
ROC-AUC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
Max Claim Probability .614 .751 .298
@@ C.1 Faithfulness Evaluation on Long-Form QA
Table 9: Llama3b results on long-form QA benchmark with faithfulness target.
C.2 Factuality under faithful/unfaithful conditions
Method Llama-3b
MeanRank $\downarrow$ MeanValue $\uparrow$
ROC-AUC PR-AUC PRR ROC-AUC PR-AUC PRR
Max Sequence Probability 7.00 7.00 7.50 .754 .518 .454
CCP 9.00 7.75 8.75 .742 .512 .434
Max Token Entropy 5.25 5.25 5.00 .767 .540 .472
P(True) 15.25 15.25 15.25 .486 .252 -.001
Lexical Similarity 7.00 7.00 7.00 .758 .500 .457
Degree Matrix 4.50 4.00 4.25 .770 .549 .488
Eccentricity 5.50 5.25 4.75 .768 .547 .482
Sum of Eigenvalues 5.50 6.50 6.25 .767 .538 .476
Number of Semantic Sets 13.00 12.50 12.50 .625 .383 .232
Semantic Entropy 2.50 3.00 2.50 .781 .562 .510
Perplexity 8.25 7.50 8.00 .741 .509 .426
TokenSAR 9.25 9.25 9.75 .733 .493 .408
SentenceSAR 5.00 6.75 5.50 .766 .518 .473
MaxNLI 15.50 15.25 15.50 .454 .246 -.067
AlignScore 13.25 13.50 13.50 .606 .321 .170
Parametric Knowledge 10.25 10.25 10.00 .657 .413 .295
(a) Only claims with AlignScore > 0.5
Method Llama-3b
MeanRank $\downarrow$ MeanValue $\uparrow$
ROC-AUC PR-AUC PRR ROC-AUC PR-AUC PRR
Max Sequence Probability 8.50 8.00 8.50 .752 .648 .446
CCP 7.75 9.25 8.00 .741 .631 .445
Max Token Entropy 6.75 5.00 6.75 .755 .673 .462
P(True) 15.50 15.50 15.25 .474 .405 -.016
Lexical Similarity 6.25 8.00 7.00 .767 .662 .469
Degree Matrix 3.00 2.75 2.50 .796 .728 .551
Eccentricity 3.50 3.50 4.25 .793 .709 .527
Sum of Eigenvalues 1.50 2.50 2.00 .807 .735 .560
Number of Semantic Sets 11.50 9.50 10.00 .693 .605 .377
Semantic Entropy 3.75 4.25 5.25 .782 .689 .502
Perplexity 9.25 8.75 9.25 .725 .634 .411
TokenSAR 11.00 10.50 10.50 .710 .616 .390
SentenceSAR 6.25 6.75 5.50 .770 .667 .473
MaxNLI 15.00 14.75 14.75 .490 .443 .043
AlignScore 14.00 14.00 14.25 .555 .488 .142
Parametric Knowledge 12.50 13.00 12.25 .602 .512 .230
(b) Only claims with AlignScore < 0.5
Table 10: Results averaged across 4 QA datasets for Llama 3B considering only claims with high and low AlignScore.
Method Llama3b
ROC-AUC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
MCP .538 .115 .028
PTrue .463 .112 .002
@@ C.2 Factuality under faithful/unfaithful conditions
Table 11: Results on long-form QA benchmark with factuality target (only unfaithful claims).
Tables 10 present the results for high and low AlignScore, respectively. The results indicate that Semantic Entropy is the top-performing method for faithful claims, and sum of Eigenvalues of Graph Laplacian is the top-performing method for unfaithful claims.
Table 11 presents the results of all unsupervised methods with only unfaithful claims left in the test. Parametric Knowledge outperforms other methods with respect to ROC-AUC and PRR metrics.
@@ Appendix D Ablation studies
D.1 franq with different choice of $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$
(a) PRR on long-form QA dataset.
(b) PRR on short-form QA benchmark (mean value across 4 datasets).
Figure 11 presents the PRR of franq condition-calibrated, for different choices of $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ methods. On the long-form dataset (Figure 11(a)), the highest PRR is achieved using MaxNLI as $\text{UQ}_{\text{faith}}$ and Parametric Knowledge as $\text{UQ}_{\text{unfaith}}$. The results indicate that on the long-form dataset, the selection of $\text{UQ}_{\text{faith}}$ has a minimal impact, while opting for Parametric Knowledge as the $\text{UQ}_{\text{unfaith}}$ method is essential.
The short-form QA results (see Figure 11(b)) show that many combinations of $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ achieve PR-AUC scores close to the top, indicating flexibility in selecting these components. The pair chosen for short-form QA experiments (Semantic Entropy and Sum of Eigenvalues) achieves a PR-AUC of 0.541, which is close to the highest score of 0.553 obtained by Max Token Entropy and Sum of Eigenvalues.
D.2 franq with alternative faithfulness estimations
Method Llama3b, long-form QA
Llama3b, short-form QA
(mean values across 4 datasets)
ROC-AUC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$ ROC-AUC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
@@
Table 12: Comparison of franq performance on Llama3b benchmarks, when using AlignScore with and without threshold.
Method Llama3b, long-form QA
ROC-AUC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
@@
franq condition-calibrated, faithfulness-calibrated .587 .124 .112
Table 13: Comparison of franq performance on Llama3b long-form QA benchmark, when applying calibration for faithfulness estimator, AlignScore.
D.3 Impact of train size on franq
(a) Long-form QA, Llama 3B
(b) Short-form QA, Llama 3B
@@ D.3 Impact of train size on franq
D.4 Analysis of XGBoost
{forest}
for tree= draw, rounded corners, align=center, edge=->,>=latex, parent anchor=south, child anchor=north, s sep=8mm, l sep=12mm, font= [AlignScore > 0.72 [MaxNLI > 0.39, edge label=node[midway,xshift=-6pt]true […, edge label=node[midway,left]true] […, edge label=node[midway,right]false] ] [Parametric Knowledge > exp(-22.13), edge label=node[midway,xshift=6pt]false […= -0.30, edge label=node[midway,left]true] […, edge label=node[midway,right]false] ] ]
Figure 13: Top vertices of first XGBoost tree trained on franq components (AlignScore, Parametric Knowledge and MaxNLI) for long-form QA Llama 3B behchmark.
We examine the first tree from an XGBoost model trained on franq features (AlignScore, MaxNLI, and Parametric Knowledge) for long-form QA with Llama 3B. While XGBoost uses multiple trees, the first tree often captures key decision patterns.
Figure 13 presents the first several nodes in first XGBoost tree. The root splits on AlignScore. If it’s high, the model next considers MaxNLI; if low, it turns to Parametric Knowledge. This mirrors franq’s logic: leading with faithfulness assessment with AlignScore, followed by either MaxNLI or Parametric Knowledge. The tree thus exhibits structure similar of franq’s decision process.
D.5 Calibration properties of UQ methods
We evaluate the calibration properties of all our UQ methods using the Expected Calibration Error (ECE; Guo et al. (2017)). ECE quantifies the alignment between predicted confidence scores and observed accuracy. Specifically, predictions are partitioned into 10 equally spaced confidence bins. Within each bin, we compute the average predicted confidence and compare it to the empirical accuracy. Lower ECE values indicate better-calibrated models.
Table 14 reports ECE scores for both long-form QA dataset and short-form QA benchmark using the Llama 3B model. Only UQ methods that produce confidence values within the [0, 1] interval are included, as this is a prerequisite for ECE computation. Notably, the two calibrated variants of franq achieve the best calibration performance across datasets.
@@ D.5 Calibration properties of UQ methods
Max Claim Probability .72
@@
(a) Long-form QA Llama 3B dataset.
@@
Max Sequence Probability .46
CCP .23
P(True) .71
@@
Number of Semantic Sets .48
Perplexity .17
MaxNLI .51
@@
(b) Short-form QA Llama 3B benchmark (ECE is averaged across 4 QA datasets).
Table 14: Expected Calibration Error (ECE) for all tested UQ methods with Llama 3B.
Appendix E franq Examples
In Figure 15, we demonstrate the behavior of franq using three examples from a long-form QA dataset evaluated with Llama 3B. We selected three representative claims and present their corresponding franq scores for both the uncalibrated version and condition-calibrated version. The latest uses monotonic functions $f$ and $g$, fitted via isotonic regression for MaxNLI and Parametric Knowledge methods, respectively.
(a) Faithful–True. franq correctly identifies the claim as faithful and uses MaxNLI, which detects high entailment with the third retrieved passage. This results in an appropriately high franq score.
(b) Unfaithful–True. franq accurately detects the claim’s low faithfulness and assigns its factuality score based on Parametric Knowledge, which is relatively high. In the uncalibrated version, the final score is underestimated due to the uncalibrated Parametric Knowledge score. The condition-calibrated version corrects this by assigning a calibrated score of 0.99, resulting in a correctly high factuality estimate.

### v3

@@ B.3 Automatic Annotation of Claims
For faithfulness annotation, each claim is automatically assigned one of three categories: “faithful” (the context supports the statement), “unfaithful-contra” (the context contradicts it), or “unfaithful-neutral” (the context provides neither support nor contradiction). For experimental evaluation, these labels are binarized: faithful $\rightarrow$ 1, and unfaithful-contra or unfaithful-neutral $\rightarrow$ 0, since the unfaithful-contra class constitutes less than 5% of the data.
For factuality annotation, each claim is likewise assigned one of three categories: “True” (factually correct), “False” (factually incorrect), and “unverifiable” (accuracy cannot be determined without relying on the provided context). Factuality is then binarized by retaining only verifiable claims, mapping False $\rightarrow$ 1 and True $\rightarrow$ 0. The prompt used with GPT-4o-search is shown in Figure 7.
@@ B.4 Manual Enhancement of Automatic Annotation
Manual verification of the automatic annotation was performed using reliable sources identified through Google search. To validate the automatic labels and assess class balance, we compared automatic and manual annotations on randomly selected claims: 100 for Llama 3B Instruct and 76 for Falcon 3B Base. The resulting class distributions are shown in Figure 8(a) and Figure 9(a), with corresponding faithfulness comparisons in Figure 10(a, b).
Because false and unverifiable categories are particularly difficult to assess automatically, these cases were prioritized for manual review. Enhanced annotations for the same 100 (Llama 3B Instruct) and 76 (Falcon 3B Base) claims appear in Figure 8(b) and Figure 9(b). We additionally conducted a full manual re-check of all False and Unverifiable claims, yielding 359 manually reviewed claims for Llama 3B Instruct and 240 for Falcon 3B Base.
Six student annotators contributed to the study, each spending about three hours on the task. Instructions were delivered informally through oral discussion, with no written guidelines. All annotators volunteered and received no financial compensation.
To evaluate annotation consistency, we ran an agreement analysis on the 100 Llama 3B Instruct claims, each independently reviewed by two annotators (see Table 10). The results indicate generally strong alignment across annotators, particularly for factuality, while highlighting some ambiguity in borderline cases.
@@
Table 10: Inter-annotator agreement for factuality and faithfulness annotations based on 100 claims of Llama 3B Instruct. Accuracy measures raw agreement, Cohen’s Kappa adjusts for chance agreement.
(a) Before manual enhancement of automatic annotation
(b) After manual enhancement of automatic annotation
Figure 8: Balance of classes of factuality annotations for the Llama 3B Instruct model. Each matrix is based on 100 randomly selected claims, comparing annotations produced by the model with those from human annotators.
(a) Before manual enhancement of automatic annotation
(b) After manual enhancement of automatic annotation
Figure 9: Balance of classes of factuality annotations for the Falcon 3B Base model. Each matrix is based on 76 randomly selected claims, comparing annotations produced by the model with those from human annotators.
(a) Llama 3B Instruct faithfulness classes
(b) Falcon 3B Base faithfulness classes
Figure 10: Balance of classes of faithfulness annotations for Llama 3B Instruct and Falcon 3B Base models. The matrices are based on 100 and 76 randomly selected claims, correspondingly, comparing annotations produced by the model with those from human annotators.
Appendix C Additional Faithfulness-Related Analysis
In this section, we provide additional analysis of the behavior and effectiveness of AlignScore as a faithfulness estimator within the FRANQ framework.
C.1 Faithfulness Distribution and Calibration
We examine the empirical behavior of AlignScore when used to estimate claim-level faithfulness. Figure 12 shows the distribution of AlignScore values computed between model-generated claims and their corresponding retrieved documents on the long-form QA benchmark, for two representative models.
We observe that a substantial fraction of claims receive intermediate faithfulness scores, reflecting cases where claims are only partially supported or rely on implicit inferences from the retrieved evidence. Across both models, more than 40% of claims fall within the range $[0.1,0.9]$.
AlignScore also demonstrates strong calibration with respect to gold faithfulness labels, achieving low expected calibration error (ECE =0.05=0.05). This indicates that AlignScore provides a reliable continuous estimate of faithfulness suitable for probabilistic combination in franq equation 2.
Question: How and when to harvest chestnuts? Retrieved passage (excerpt): “When to harvest chestnuts? Chestnuts don’t all ripen at once. Harvest typically spans up to five weeks, but most nuts ripen within a 10–30 day period in late August and September.” Model-generated claim: “The best time to harvest chestnuts is during the 10–30 day ripening window.” AlignScore: 0.61 Analysis: While the retrieved passage mentions a 10–30 day ripening period, it does not specify that this window constitutes the best time for harvesting. Accordingly, AlignScore assigns this claim an intermediate faithfulness score of 0.61, reflecting partial grounding.
Figure 11: Example illustrating intermediate AlignScore values arising from partial grounding between a claim and retrieved evidence.
Figure 11 provides a representative qualitative example illustrating how intermediate faithfulness values arise in practice.
Figure 12: Distribution of AlignScore-based faithfulness estimates on the long-form QA benchmark for Llama 3B Instruct and Falcon 3B Base. A substantial mass lies in the intermediate region, and low ECE values indicate good calibration.
C.2 Faithfulness Evaluation on Long-Form QA
Method Llama 3B Instruct
AUROC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
General Baselines
Max Claim Prob. .614 .751 .298
@@ C.2 Faithfulness Evaluation on Long-Form QA
(a) Results on long-form QA benchmark with faithfulness target.
Method Llama 3B Instruct
AUROC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
General Baselines
Max Claim Prob. .538 .115 .028
P(True) .463 .112 .002
@@
(b) Results on long-form QA benchmark with factuality target (only unfaithful claims).
Table 11: Additional faithfulness-related results
We next evaluate the effectiveness of AlignScore as a faithfulness estimator on the long-form QA benchmark. Table 11(a) reports performance when faithfulness is treated as the target metric. All methods follow the same experimental setup used for the factuality evaluation.
Among the compared approaches, AlignScore achieves the strongest performance across metrics, indicating its effectiveness in approximating claim-level faithfulness within the FRANQ decomposition.
C.3 Factuality Under Faithful and Unfaithful Conditions
We further analyze factuality estimation under faithful and unfaithful conditions. Table 11(b) reports results for unsupervised methods when restricting evaluation to unfaithful claims only. In this setting, methods leveraging parametric knowledge perform best, achieving the highest AUROC and PRR scores.
Tables 12 report results averaged across four QA datasets for Llama 3B Instruct, considering only claims with high and low AlignScore, respectively. For faithful claims, Semantic Entropy achieves the best performance, whereas for unfaithful claims, the sum of eigenvalues of the Graph Laplacian performs best. These results further motivate the use of different uncertainty estimators conditioned on faithfulness within FRANQ.
Method Llama 3B Instruct
AUROC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
General Baselines
Max Sequence Prob. .754 .518 .454
Mean Token Entropy .767 .540 .472
CCP .742 .512 .434
Lexical Similarity .758 .500 .457
Degree Matrix .770 .549 .488
Sum of Eigenvalues .767 .538 .476
Semantic Entropy .781 .562 .510
SentenceSAR .766 .518 .473
RAG-Specific Baselines
AlignScore .606 .321 .170
Parametric Knowledge .657 .413 .295
(a) Only claims with AlignScore > 0.5
Method Llama 3B Instruct
AUROC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
General Baselines
Max Sequence Prob. .752 .648 .446
Mean Token Entropy .755 .673 .462
CCP .741 .631 .445
Lexical Similarity .767 .662 .469
Degree Matrix .796 .728 .551
Sum of Eigenvalues .807 .735 .560
Semantic Entropy .782 .689 .502
SentenceSAR .770 .667 .473
RAG-Specific Baselines
AlignScore .555 .488 .142
Parametric Knowledge .602 .512 .230
(b) Only claims with AlignScore < 0.5
Table 12: Results averaged across 4 QA datasets for Llama 3B Instruct considering only claims with high and low AlignScore.
@@ Appendix D Ablation Studies
D.1 franq with Alternative Faithfulness Estimators
Method
Llama 3B Instruct,
long-form QA
Llama 3B Instruct
mean across 4 short-form QA
AUROC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$ AUROC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
@@
Table 13: Comparison of franq performance on Llama 3B Instruct benchmarks, when using AlignScore with and without threshold.
Method Llama 3B Instruct, long-form QA
AUROC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
@@
franq condition-, faithfulness-calibrated .587 .124 .112
Table 14: Comparison of franq performance on Llama 3B Instruct long-form QA benchmark, when applying calibration for faithfulness estimator, AlignScore.
D.2 Impact of Train Size on franq
(a) Long-form QA, Llama 3B Instruct
(b) Short-form QA, Llama 3B Instruct
@@ D.2 Impact of Train Size on franq
D.3 Analysis of XGBoost
We examine the first tree from an XGBoost model trained on franq features (AlignScore, Claim Probability, and Parametric Knowledge) for long-form QA with Llama 3B Instruct. While XGBoost uses multiple trees, the first tree often captures key decision patterns.
Figure 14 presents the first several nodes in first XGBoost tree. The root splits on AlignScore. If it’s high, the model next considers Claim Probability; if low, it turns to Parametric Knowledge. This mirrors franq’s logic: leading with faithfulness assessment with AlignScore, followed by either Claim Probability or Parametric Knowledge. The tree thus exhibits structure similar of franq’s decision process.
AlignScore > 0.72
ClaimProb > 0.39
Parametric Knowledge > exp(-22.13)
truetruefalsefalsetruefalse
Figure 14: Top vertices of first XGBoost tree trained on franq components (ClaimProb) for long-form QA Llama 3B Instruct behchmark.
D.4 Calibration Properties of UQ Methods
We evaluate the calibration properties of all our UQ methods using the Expected Calibration Error (ECE; Guo et al., 2017). ECE quantifies the alignment between predicted confidence scores and observed accuracy. Specifically, predictions are partitioned into 10 equally spaced confidence bins. Within each bin, we compute the average predicted confidence and compare it to the empirical accuracy. Lower ECE values indicate better-calibrated models.
Table 15 reports ECE scores for both long-form QA dataset and short-form QA benchmark using the Llama 3B Instruct model. Only UQ methods that produce confidence values within the [0, 1] interval are included, as this is a prerequisite for ECE computation. Notably, the two calibrated variants of franq achieve the best calibration performance across datasets.
@@ D.4 Calibration Properties of UQ Methods
General Baselines
Max Claim Prob. .72
@@
(a) Long-form QA Llama 3B Instruct dataset.
@@
General Baselines
Max Sequence Prob. .46
@@
CCP .23
RAG-Specific Baselines
@@
(b) Short-form QA Llama 3B Instruct benchmark (ECE is averaged across 4 QA datasets).
Table 15: Expected Calibration Error (ECE) for all tested UQ methods with Llama 3B Instruct.
Method Inference Mean Runtime Training Time Model Size
Max Claim Probability <0.1<0.1 s
P(True) 1.3 s
Perplexity <0.1<0.1 s
Max Token Entropy <0.1<0.1 s
CCP 1.7 s
AlignScore 0.5 s
Parametric Knowledge 1.6 s
XGBoost (all UQ features) 1.9 s 0.60 s 10 kB
XGBoost (FRANQ features) 1.7 s 0.12 s 14 kB
FRANQ (no calibration) 1.7 s
FRANQ (calibrated) 1.7 s 0.57 s 312 B
FRANQ (condition-calibrated) 1.7 s 0.58 s 244 B
Table 16: Inference runtime, training time, and model size for uncertainty estimators and FRANQ variants, measured on Llama 3B Instruct for short-form QA. All runtimes represent overhead beyond base LLM generation.
Appendix E Runtime Analysis
In Table 16 we present the runtime overhead, training cost, and model size of franq’s uncertainty components. All runtimes are measured as incremental cost beyond a completed LLM forward pass using Llama 3B Instruct on short-form NQ dataset (average generation time: 2.2 s per instance).
Overall, franq incurs modest overhead, while isotonic calibration adds negligible cost and yields compact models (hundreds of bytes), supporting production deployment.
Appendix F Resource and Expenses
A full data-generation and UQ-baseline evaluation run required about 8 days of compute on an NVIDIA V100 32GB GPU for long-form QA, while short-form QA needed under one day. The OpenAI API was used for claim splitting, matching, and annotation, costing roughly \$100 per model run (Llama 3B Instruct). Human annotation involved six student annotators, each contributing about three hours of work.
Appendix G franq Examples
In Figure 15, we demonstrate the behavior of franq using three examples from a long-form QA dataset evaluated with Llama 3B Instruct. We selected three representative claims and present their corresponding franq scores for both the uncalibrated version and condition-calibrated version. The latest uses monotonic functions $f$ and $g$, fitted via isotonic regression for Claim Probability and Parametric Knowledge methods, respectively.
(a) Faithful–True. franq correctly identifies the claim as faithful and uses Claim Probability, which detects high entailment with the third retrieved passage. This results in an appropriately high franq score.
(b) Unfaithful–True. franq accurately detects the claim’s low faithfulness and assigns its factuality score based on Parametric Knowledge, which is relatively high. In the uncalibrated version, the final score is underestimated due to the uncalibrated Parametric Knowledge score. The condition-calibrated version corrects this by assigning a calibrated score of 0.85, resulting in a correctly high factuality estimate.